# NeuroNAS: A Framework for Energy-Efficient Neuromorphic Compute-in-Memory Systems using Hardware-Aware Spiking Neural Architecture Search

### v2

@@ II-B CIM Hardware Architecture and SNN Mapping Strategy
  • A similar partitioning is applied to ifms over the $P$$\times$$P$$\times$$D$ block.
Figure 7: SNN partitioning for mapping input feature maps (ifms) and weights (wghs) on the CIM crossbars; adapted from studies in [18]. The dimension of ifms is denoted as $H$$\times$$H$$\times$$D$, while wghs is denoted as $P$$\times$$P$$\times$$D$$\times$$F$.
Furthermore, this mapping strategy considers a design choice that a network layer can be mapped on multiple CIM tiles, but multiple network layers cannot be mapped in one CIM tile [18].
@@ III The NeuroNAS Framework
We propose NeuroNAS framework to address the research challenges, whose key steps are discussed below (see overview in Fig. 5).
@@ III-A Optimization of SNN Operations
  • Scenario-2: We investigate the impact of other possible operations with different kernel sizes (i.e., 5x5Conv and 7x7Conv) on accuracy and memory cost, to explore their potentials as alternative solutions. To do this, we replace the operation that has the role of extracting features from pre-define ones (i.e., 3x3Conv) with the investigated operation (either 5x5Conv or 7x7Conv).
All these scenarios consider a cell-based SNN architecture shown in Fig. 4 and the CIFAR-100 workload. Experimental results are shown in Fig. 8, from which we make the following key observations.
@@
    Eliminating Zeroize, 1x1Conv, and 3x3AvgPool from the search space does not significantly degrade th accuracy as compared to the baseline. Hence, these operations can be removed from the search space. If model compression is needed for reducing the memory size, we can keep 3x3AvgPool in the search space.
@@
    Eliminating SkipCon from the search space slightly reduces accuracy, since SkipCon is useful for providing feature maps from previous layer to preserve important information. Similar results are also observed in [33]. Hence, SkipCon should be kept in the search space to maintain the learning quality.
@@
    Eliminating 3x3Conv from the search space significantly degrades accuracy, since the main role of 3x3Conv is to extract important features from input samples. Hence, 3x3Conv should be kept in the search space to maintain the learning quality.
@@
These observations show the order of significance of the operations for achieving high accuracy with low memory cost, from the highest to the lowest: (1) 3x3Conv, (2) SkipCon, (3) 3x3AvgPool, then (4) 5x5Conv, 7x7Conv, Zeroize, and 1x1Conv. This information is then leveraged in Sec. III-B1.
Figure 8: (a) Results of eliminating different operations from the search space of the cell architecture. (b) Results of investigating the impact of 5x5Conv and 7x7Conv. Here, Eval_3x3, Eval_5x5, and Eval_7x7 refer to then accuracy when evaluating the impact of 3x3Conv, 5x5Conv, and 7x7Conv, respectively.
@@ III-B Development of SNN Architecture
This step aims to build an SNN architecture that facilitates effective learning. Toward this, it considers two parts: (1) development of the neural cell architecture, and (2) development of the SNN architecture.
@@ III-B1 Neural Cell Architecture
Our discussion in Sec. III-A show that we can optimize the search space of cell-based NAS strategy for maintaining/improving the accuracy while keeping the memory small. To do this, we select a few operations that have positive impact on accuracy with low memory cost to be considered in the search space. Specifically, we consider SkipCon, 3x3Conv, and 3x3AvgPool in the search space, while keeping 4 nodes of DAG in the neural cell.
@@ III-B2 SNN Architecture
The CIM platforms typically store the SNN weights in NVM devices (e.g., RRAM) using a fixed-point format due to its efficient implementation [18]. How each weight value stored in CIM hardware depends on the precision of NVM devices and SNN weights. For instance, a multi-bit weight value may be stored in several single-level-cell (SLC) devices, or in a single multi-level-cell (MLC) device. This indicates that SNNs should be developed and optimized considering the CIM parameters (e.g., NVM device precision) to ensure the deployability of SNNs. However, simply performing post-training quantization (PTQ) with fixed-point format may significantly degrade the accuracy; see 5 in Fig. 9(a). To address this, we propose a Quantization-aware Fitness Evaluation (QaFE) policy for quantizing the network candidates during NAS, thereby compressing the network size while improving the fitness function. Our QaFE policy employs the following key steps; see Fig. 9(b).
@@
  • Weight quantization is then performed for the investigated network based on the precision level and the rounding scheme.
  • Afterward, mini-batch samples $S$ are fed to the network to calculate the Hamming distances between different samples, build a matrix $\textbf{M}_{H}$, and calculate the fitness score $F$.
In this manner, the fitness function is enhanced to better reflect the representation capability of the quantized network rather than the fitness score from the original network; see 6 in Fig. 9(a).
@@ III-C A Hardware-aware Search Algorithm
  • Each cell has an individual search to determine its internal architecture by exploring possible candidates from different combinations of architectures from multiple cells: Cell-A and Cell-B.
  • First, we explore possible architectures inside the Cell-A (Alg. 1: line 3-5). Here, the Cell-B architecture is initially set the same as Cell-A (Alg. 1: line 6). Then, we build an SNN based on Cell-A and Cell-B, and quantize it (Alg. 1: line 7-8).
  • Afterward, we evaluate the SNN memory cost considering the number of SNN weight parameters (Alg. 1: line 9-10). If memory constraint is met, then we perform the fitness evaluation (Alg. 1: line 11). If the fitness score is higher than the saved one, we record the SNN architecture as a solution candidate (Alg. 1: line 12-16).
  • Second, we explore possible architectures inside the Cell-B (Alg. 1: line 3 and 17-18). Then, we build an SNN based on the states of Cell-A and Cell-B, and quantize it (Alg. 1: line 19-20).
  • Afterward, we evaluate the candidate for its memory, area, latency, and energy costs (Alg. 1: line 21-25). If all constraints are met, we proceed with the fitness function evaluation (Alg. 1: line 26). If the fitness score is higher than the saved one, then we record the SNN architecture as a solution candidate (Alg. 1: line 27-30).
  • After all steps are finished, we consider the saved SNN architecture as the solution of NAS (Alg. 1: line 31) which is ready to train.
To enable quick and accurate area, latency, and energy calculations in the NAS, we also propose adjustments considering the precision of NVM device ($bit_{D}$) and SNN weights ($bit_{W}$). Specifically, we propose to employ $\lceil bit_{W}/bit_{D}\rceil$ as an adjustment factor of the corresponding calculations for allocating the crossbar area and calculating the energy consumption, and which also affect the processing latency.
@@
0:[1] Number of cell operations ($P$): SkipCon=0, 3x3Conv=1, 3x3AvgPool=2; [2] Design requirements and constraints: memory ($const_{mem}$), area ($const_{mem}$), latency ($const_{lat}$), energy consumption ($const_{eng}$);
@@
11:             $score=$ fitness_eval$(arch_{q})$;
@@
29:             $score=$ fitness_eval$(arch_{q})$;
@@
31:                $net\_arch=arch_{q}$;
32:                $score_{h}=score$;
33:                $saved\_cell_{B}=cell_{B}$;
@@ IV Evaluation Methodology
Software Evaluation: To evaluate our NeuroNAS framework, we develop a PyTorch implementation based on the SpikingJelly [38], then run it on the Nvidia GeForce RTX A6000 GPU machines. We train the network using a surrogate gradient learning [39][40] with 200 epochs. To show the generality of our solution, we consider the CIFAR-10, CIFAR-100, and TinyImageNet-200 which represent simple, medium, and complex datasets, respectively. For the comparison partner, we consider the state-of-the-art SNASNet with its default settings (e.g., 5000x random search iterations) [2]. We also assume 32bit fixed-point (FxP32) for SNASNet implementation on CIM hardware, as there is no quantization in SNASNet. Outputs of this evaluation are the accuracy, memory cost, and searching time.
Hardware Evaluation: To evaluate the area, latency, and energy consumption, we employ the energy-latency-area (ELA) simulator while considering the SpikeFlow CIM hardware architecture with real measurements-based RRAM device values from the studies in [18]. The circuit and device parameters are shown in Table I.
TABLE I: The CIM circuit and NVM (RRAM) device parameters considered in the ELA simulator; adapted from studies in [18].
@@ V-A Maintaining High Accuracy
Fig. 11 shows the experimental results of the accuracy. The state-of-the-art SNASNet achieves 90.79%, 67.18%, and 53% accuracy for the CIFAR-10, CIFAR-100, and TinyImageNet-200, respectively. These are obtained through the network generation in SNASNet, which simply relies on the random search iteration. Meanwhile, our NeuroNAS can achieve comparable accuracy to the state-of-the-art with smaller precision levels (i.e., with at least 8bits), as shown by 1 for CIFAR-10, 2 for CIFAR-100, and 3 for TinyImageNet-200. For CIFAR-10, NeuroNAS achieves accuracy of 91.44% (FxP16), 90.53% (FxP14), 88.92% (FxP12), 88.27% (FxP10), and 83.44% (FxP8). For CIFAR-100, NeuroNAS achieves accuracy of 63.90% (FxP16), 60.74% (FxP14), 59.87% (FxP12), 59.83% (FxP10), and 57.59% (FxP8). Meanwhile, for TinyImageNet-200, NeuroNAS achieves accuracy of 53% (FxP16), 53% (FxP14), 57% (FxP12), 59% (FxP10), and 53% (FxP8). These results indicate that, our NeuroNAS can still preserve the high accuracy as compared to the state-of-the-art across different datasets while employing smaller precision levels. Moreover, in the TinyImageNet-200 case, NeuroNAS can consistently achieve similar or even higher accuracy to the state-of-the-art. The reasons are the following: (1) NeuroNAS only employs cell operations that have high significance and positive impact on the accuracy, and (2) NeuroNAS constructs the network by considering its fitness score under quantized weights, hence enabling more accurate evaluation of network candidates and search toward the desired SNN architecture.
Figure 11: Experimental results of the test accuracy achieved by the SNASNet and our NeuroNAS across training epochs and different precision levels (i.e., FxP16, FxP14, FxP12, FxP10, FxP8, FxP6, and FxP14) for (a) CIFAR-10, (b) CIFAR-100, and (c) TinyImageNet-200 workloads.
Figure 12: Experimental results of the hardware area, processing latency, and energy consumption achieved by the SNASNet and our NeuroNAS across different precision levels (i.e., FxP16, FxP14, FxP12, FxP10, FxP8, FxP6, and FxP14) for (a) CIFAR-10, (b) CIFAR-100, and (c) TinyImageNet-200 workloads.
Figure 13: Experimental results of the searching time achieved by the SNASNet and our NeuroNAS across different precision levels (i.e., FxP16, FxP14, FxP12, FxP10, FxP8, FxP6, and FxP14) for (a) CIFAR-10, (b) CIFAR-100, and (c) TinyImageNet-200 workloads.
@@ V-B Ensuring Hardware-aware SNN Generation
Fig. 12 shows the experimental results for hardware-related evaluation, i.e, memory, area, latency, and energy consumption.
Memory: Both SNASNet and our NeuroNAS satisfy the memory constraint. SNASNet incurs 1.28M weight parameters for CIFAR-10, and 1.93M for CIFAR-100 and TinyImageNet-200. Meanwhile, NeuroNAS incurs  [^8] K-4.7M weight parameters for CIFAR-10, 1.77M-4.72M for CIFAR-100, and 1.77M-3.83M for TinyImageNet-200 across different precisions (i.e., FxP4-FxP16). These are obtained as NeuroNAS incorporates the memory constraint in its NAS algorithm.
Area: SNASNet generates networks that occupy area of 634mm2 for CIFAR-10 and CIFAR-100, and 755mm2 for TinyImageNet-200; see 4. Since SNASNet does not consider the area constraint, the hardware area of the generated network is not under control during its NAS process. Moreover, in each neural cell, there are multiple convolutional (CONV) layers which cannot be mapped in the same CIM tile due to the mapping strategy. This mapping forces the CIM processing to utilize more resources (e.g., CIM tiles). Meanwhile, our NeuroNAS generates networks that occupy area of 26mm2-187mm2 for CIFAR-10 (70%-96% savings from SNASNet), 25mm2-209mm2 for CIFAR-100 (67%-96% savings from SNASNet), and 25mm2-166mm2 for TinyImageNet-200 (78%-96% savings from SNASNet) across different precision levels; see 5. These indicate that NeuroNAS can ensure the generated network to meet the area constraint, which also leads to smaller area costs compared to SNASNet. The reason is that, NeuroNAS includes the area requirement in its NAS process, hence making it possible to develop both networks and CIM hardware within a custom area budget.
Latency: SNASNet generates networks that have latency of 468ms for both CIFAR-10 and CIFAR-100, and 545ms for TinyImageNet-200, which does not meet the latency constraint; see 4. The reason is that, SNASNet does not consider the latency budget in its NAS process. Meanwhile, our NeuroNAS generates networks that have latency of 340ms-494ms for CIFAR-10, 315ms-494mm for CIFAR-100 and TinyImageNet-200 across different precisions; see 5. These show that NeuroNAS can ensure the generated network to fulfill the latency constraint, since it considers the latency information in its NAS process, hence making it possible to develop both networks and CIM hardware within a custom latency budget.

### v3

@@ II-C IMC Hardware Architecture and SNN Mapping Strategy
  • A similar data partitioning strategy is applied to ifms over the $P$$\times$$P$$\times$$D$ block.
Furthermore, this mapping strategy considers a design choice that a network layer can be mapped on multiple tiles, but multiple network layers cannot be mapped in one tile [4].
@@ III The NeuroNAS Framework
We propose NeuroNAS framework, whose key steps are discussed below (see overview in Fig. 7).
@@ III-A Optimization of SNN Operations
  • Scenario-2: We investigate the impact of other possible operations with different kernel sizes (i.e., 5x5Conv and 7x7Conv) on accuracy and memory cost, to explore their potentials as alternative solutions. Here, we replace the operation that has the role of extracting features from pre-define ones (i.e., 3x3Conv) with the investigated operation (either 5x5Conv or 7x7Conv).
All these scenarios consider a cell-based SNN illustrated in Fig. 4 and CIFAR-100. Experimental results are shown in Fig. 8, from which we make the following key observations.
@@
    Eliminating Zeroize, 1x1Conv, and 3x3AvgPool from the search space does not significantly degrade accuracy compared to the baseline. Hence, these operations can be removed from the search space. If model compression is needed for reducing the memory size, we can keep 3x3AvgPool in the search space.
@@
    Eliminating SkipCon from the search space slightly reduces accuracy, since SkipCon is useful for providing feature maps from previous layer to preserve important information. Similar results are also observed in [14]. Hence, SkipCon should be kept in the search space.
@@
    Eliminating 3x3Conv from the search space significantly degrades accuracy, as the main role of 3x3Conv is to extract important features from input samples. Hence, 3x3Conv should be kept in the search space to maintain the learning quality.
@@
These observations show the order of significance of the operations for achieving high accuracy with low memory cost, from the highest to the lowest: (1) 3x3Conv, (2) SkipCon, (3) 3x3AvgPool, then (4) 5x5Conv, 7x7Conv, Zeroize, and 1x1Conv. This insight is then leveraged in Sec. III-B1.
Figure 8: (a) Results of eliminating different operations from the search space of the cell architecture. (b) Results of investigating the impact of 5x5Conv and 7x7Conv. Eval_3x3, Eval_5x5, and Eval_7x7 refer to then accuracy when evaluating the impact of 3x3Conv, 5x5Conv, and 7x7Conv, respectively.
@@ III-B Development of SNN Architecture
This step aims to build an SNN architecture that facilitates effective learning. It considers two parts: developments of (1) the neural cell architecture, and (2) the SNN architecture.
@@ III-B1 Neural Cell Architecture
Based on our analysis in Sec. III-A, we can optimize the search space of cell-based NAS strategy for maintaining/improving the accuracy while keeping the memory small. To do this, we select a few operations that have positive impact on accuracy with low memory cost to be considered in the search space. Specifically, we consider SkipCon, 3x3Conv, and 3x3AvgPool in the search space, while keeping 4 nodes of DAG in the neural cell.
@@ III-B2 SNN Architecture
The IMC platforms typically store SNN weights in NVM devices (e.g., RRAM) using a fixed-point format due to its efficient implementation [4]. Consequently, how each weight value stored in the IMC hardware depends on the NVM device precision. For instance, a multi-bit weight value may be stored in several single-level-cell (SLC) devices, or in a single multi-level-cell (MLC) device. This indicates that SNNs should be developed and optimized considering the IMC parameters (e.g., NVM device precision) to ensure the deployability of SNNs. However, simply performing post-training quantization (PTQ) with fixed-point format may significantly degrade the accuracy; see 5 in Fig. 9(a). To address this, we propose a Quantization-aware Fitness Evaluation (QaFE) policy for quantizing the network candidates during NAS, thereby compressing the network size while improving the fitness function. Our QaFE policy employs the following key steps; see Fig. 9(b).
@@
  • Weight quantization is then performed on the network based on the precision level and the rounding scheme.
  • Then, mini-batch samples $S$ are fed to the network to calculate the Hamming distances between different samples, build a matrix $\textbf{M}_{H}$, and calculate the fitness score $F$.
This way, the fitness function is enhanced to better reflect the representation capability of the quantized network than the fitness score from the original network; see 6 in Fig. 9(a).
@@ III-C A Hardware-aware Search Algorithm
0:  (1) Number of neural cell operations ($P$): SkipCon=0, 3x3Conv=1, 3x3AvgPool=2; (2) Application requirements & constraints: memory ($const_{mem}$), area ($const_{mem}$), latency ($const_{lat}$), energy consumption ($const_{eng}$);
@@
11:              $score=$ fitness_eval$(arch_{q})$; // evaluate the fitness score
@@
29:              $score=$ fitness_eval$(arch_{q})$; // evaluate the fitness score
@@
31:                 $score_{h}=score$;
32:                 $saved\_cell_{B}=cell_{B}$;
33:                 $net\_arch=arch_{q}$;
@@
  • First, we explore possible architectures inside the Cell-A (lines 3-5). Here, the Cell-B architecture is initially set the same as Cell-A (line 6). Then, we build an SNN based on Cell-A and Cell-B, and quantize it (lines 7-8).
  • We evaluate the SNN memory cost, i.e, number of SNN weight parameters (lines 9-10). If memory constraint is met, then we perform the fitness evaluation (line 11). If the fitness score is higher than the saved one, we record the SNN architecture as a solution candidate (lines 12-16).
  • Second, we explore possible architectures inside the Cell-B (lines 3, 20-21). Then, we build an SNN based on the states of Cell-A and Cell-B, and quantize it (lines 22-23).
  • We evaluate the candidate for its memory, area, latency, and energy costs (lines 24-28). If all constraints are met, we perform the fitness function evaluation (line 29). If the fitness score is higher than the saved one, then we record the SNN architecture as a solution candidate (lines 30-33).
  • After all steps are finished, we consider the saved SNN architecture as the NAS solution which is ready to train.
Furthermore, to enable quick and accurate area, latency, and energy calculations in the NAS, we also propose adjustments considering the precision of NVM device ($bit_{D}$) and SNN weights ($bit_{W}$). Specifically, we employ $\lceil bit_{W}/bit_{D}\rceil$ as an adjustment factor of the corresponding calculations for allocating the crossbar area and calculating the energy consumption, and hence the processing latency.
@@ IV Evaluation Methodology
Software Evaluation: We realize the NeuroNAS framework using PyTorch implementation based on the SpikingJelly [30], then run it on the Nvidia GeForce RTX A6000 Multi-GPU machines. We train the network using a surrogate gradient learning with 200 epochs. To show the generality of our framework, we use CIFAR-10, CIFAR-100, and TinyImageNet-200 which represent simple, medium, and complex datasets, respectively. For the comparison partner, we consider the state-of-the-art SNASNet with its default settings (e.g., 5000x random search iterations) [15]. We use a 32-bit fixed-point (FxP32) for SNASNet implementation on IMC hardware, since there is no quantization in SNASNet. The outputs include accuracy, memory, and searching time.
Hardware Evaluation: To evaluate area, processing latency, and energy consumption, we employ the state-of-the-art energy-latency-area (ELA) simulator while considering the SpikeFlow IMC hardware architecture with real measurements-based RRAM device values from [4]. The circuit and device parameters are shown in Table I.
TABLE I: The IMC circuit and NVM (RRAM) device parameters considered in the ELA simulator; based on studies in [4].
@@ V Results and Discussion
Figure 11: Results of the test accuracy achieved by the SNASNet and our NeuroNAS across training epochs and different precision levels (i.e., FxP16, FxP14, FxP12, FxP10, FxP8, FxP6, and FxP14) and different datasets: (a) CIFAR-10, (b) CIFAR-100, and (c) TinyImageNet-200.
Figure 12: Results of the hardware area, processing latency, and energy consumption achieved by the SNASNet and our NeuroNAS across different precision levels (i.e., FxP16, FxP14, FxP12, FxP10, FxP8, FxP6, and FxP14) and different datasets: (a) CIFAR-10, (b) CIFAR-100, and (c) TinyImageNet-200.
Figure 13: Results of the searching time achieved by the SNASNet and our NeuroNAS across different precision levels (i.e., FxP16, FxP14, FxP12, FxP10, FxP8, FxP6, and FxP14) and different datasets: (a) CIFAR-10, (b) CIFAR-100, and (c) TinyImageNet-200.
@@ V-A Maintaining High Accuracy
Fig. 11 provides the experimental results for accuracy. Here, the state-of-the-art (i.e., SNASNet) achieves 90.79%, 67.18%, and 53% accuracy for CIFAR-10, CIFAR-100, and TinyImageNet-200, respectively. These are obtained through the network generation in SNASNet, which simply relies on the random search iteration. Meanwhile, our NeuroNAS can achieve comparable accuracy to the state-of-the-art with smaller precision levels (i.e., with at least 8bits), as shown by 1 for CIFAR-10, 2 for CIFAR-100, and 3 for TinyImageNet-200. For CIFAR-10, NeuroNAS achieves accuracy of 91.44% (FxP16), 90.53% (FxP14), 88.92% (FxP12), 88.27% (FxP10), and 83.44% (FxP8). For CIFAR-100, NeuroNAS achieves accuracy of 63.90% (FxP16), 60.74% (FxP14), 59.87% (FxP12), 59.83% (FxP10), and 57.59% (FxP8). Meanwhile, for TinyImageNet-200, NeuroNAS achieves accuracy of 53% (FxP16), 53% (FxP14), 57% (FxP12), 59% (FxP10), and 53% (FxP8). These results show that, our NeuroNAS preserves high accuracy comparable to that of the state-of-the-art across different datasets while employing smaller precision levels. Moreover, in the TinyImageNet-200 case, NeuroNAS can consistently achieve similar or even higher accuracy to the state-of-the-art. The reasons are the following: (1) NeuroNAS only employs cell operations that have high significance and positive impact on the accuracy, and (2) NeuroNAS constructs the network by considering its fitness score under quantized weights, hence enabling more accurate evaluation of network candidates and search toward the desired SNN architecture.
@@ V-B Ensuring Hardware-aware SNN Generation
Fig. 12 provides the experimental results for memory, area, latency, and energy consumption.
Memory: Both SNASNet and our NeuroNAS satisfy the memory constraint. SNASNet incurs 1.28M weight parameters for CIFAR-10, and 1.93M for CIFAR-100 and TinyImageNet-200. Meanwhile, NeuroNAS incurs  [^8] K-4.7M weight parameters for CIFAR-10, 1.77M-4.72M for CIFAR-100, and 1.77M-3.83M for TinyImageNet-200 across different precisions (i.e., FxP4-FxP16). These are obtained as NeuroNAS incorporates the memory constraint in its NAS.
Area: SNASNet generates networks that occupy area of 634mm2 for CIFAR-10 and CIFAR-100, and 755mm2 for TinyImageNet-200; see 4. Since SNASNet does not consider the area constraint, the hardware area of the generated network is not under control during its NAS process. Moreover, in each neural cell, there are multiple convolutional (CONV) layers which cannot be mapped in the same IMC tile due to the mapping strategy. This mapping forces the IMC processing to utilize more resources (e.g., IMC tiles). Meanwhile, our NeuroNAS generates networks that occupy area of 26mm2-187mm2 for CIFAR-10 (70%-96% savings from SNASNet), 25mm2-209mm2 for CIFAR-100 (67%-96% savings from SNASNet), and 25mm2-166mm2 for TinyImageNet-200 (78%-96% savings from SNASNet) across different precision levels; see 5. These indicate that NeuroNAS can ensure the generated network to meet the area constraint, which also leads to smaller area costs compared to SNASNet. The reason is that, NeuroNAS includes the area constraint in its NAS, making it possible to develop both networks and IMC hardware within a custom area budget.
Latency: SNASNet generates networks that have latency of 468ms for both CIFAR-10 and CIFAR-100, and 545ms for TinyImageNet-200, which does not meet the latency constraint; see 4. The reason is that, SNASNet does not consider the latency budget in its NAS process. Meanwhile, our NeuroNAS generates networks that have latency of 340ms-494ms for CIFAR-10, 315ms-494mm for CIFAR-100 and TinyImageNet-200 across different precisions; see 5. These show that NeuroNAS can ensure the generated network to fulfill the latency constraint, since it considers the latency information in its NAS, making it possible to develop both networks and IMC hardware within a custom latency budget.